# Physics-informed Neural Networks for Heterogeneous Poroelastic Media

### v1

@@ Abstract
This study introduces a novel physics-informed neural networks (PINNs) framework designed to model coupled-field problems specifically tailored for heterogeneous poroelastic media. Firstly, a composite neural network is developed where distinct neural networks are dedicated to predicting displacement and pressure variables for each material, employing identical activation functions but trained separately across all other parameters. Secondly, we handle the challenges of heterogeneous material interfaces by the Interface-PINNs (I-PINNs) framework, where different activation functions across any material interface are prescribed to ensure that the discontinuities in solution fields and gradients are accurately captured. We compare the modified PINNs framework with the conventional approach on two one-dimensional benchmark examples for poroelasticity in heterogeneous media. Furthermore, we assess a single neural network architecture, comparing it against the composite neural network proposed in this work. These examples show that the proposed framework demonstrates superior approximation accuracy in both displacements and pressures, and better convergence behavior.
@@ 1 Introduction
Various geothermal and subsurface energy applications are governed by the dynamics of fluid flow within poroelastic media. This encompasses activities like hydrocarbon retrieval from oil and gas reservoirs [1], geological storage of carbon dioxide [2], and harnessing energy from subsurface thermal systems [3]. In reservoir engineering [4, 5], these interactions are key in determining the state of subsurface fluids and gases, the performance of production wells, and the deformation of the porous subsurface media. In geothermal systems, heat below the ground is transported to the surface through fluid flow, requiring a thorough thermo-hydro-poromechanical analysis for system optimization. Furthermore, the inherent heterogeneity characterizing these subsurface systems adds a layer of complexity. Heterogeneities may stem from the natural stratification of the earth’s crust, the presence of naturally occurring faults, or induced variations due to processes such as hydraulic fracturing [6]. The diverse nature of these systems can significantly affect the behavior and movement of fluids below the surface. Therefore, a thorough investigation and comprehension of subsurface heterogeneity, coupled with the intricate thermo-hydro-poromechanical processes, emerge as imperative aspects of studying geothermal and subsurface applications.
Computational modeling is often necessary to supplement theoretical and experimental investigations to gain fundamental insight into the coupled physical processes of flow and deformation. On the one hand, compared to a fully theoretical investigation, a computational model can relax simplifying assumptions on problem geometry, presence of material interfaces, and defects. On the other hand, compared to an experimental study, computational models are relatively inexpensive and provide a virtual laboratory to conduct numerical experiments at a fraction of the cost.
@@
Computational models of these thermo-hydro-poromechanical systems can be broadly categorized as either physics-based or data-driven. Physics-based models involve the solution of coupled partial differential equations (PDEs). These coupled PDEs are solved numerically, often employing methods like the finite element method [7, 8, 9, 10], finite volume method [11], or finite difference method [12], that provide approximate solutions to the PDEs on an underlying grid. However, generating grids that conform to subsurface features (such as material interfaces, faults, and fractures) is a time-consuming pre-processing step. For instance, in several reservoir engineering applications, the cost of grid generation is comparable to (or exceeds) the cost of analysis [13, 14, 15]. This makes computational modeling intractable for realistic field-scale applications. By contrast, data-driven models employ machine learning (ML), particularly artificial neural networks (ANNs), to predict the coupled behavior. However, their accuracy relies on the availability and quality of training data, which can be sparse in subsurface systems due to economic constraints. As such, the applicability of purely data-driven models in poroelasticity is also severely limited.
@@
Physics-informed neural networks (PINNs), a recently developed hybrid approach [16], combines aspects of both physics-based and data-driven modeling. On the one hand, PINNs are not grid-based and thereby eliminate the need for conforming meshes. On the other hand, PINNs leverage fundamental physical laws including governing PDEs, initial and boundary conditions, and data, to construct a loss functional and thereby do not require extensive training datasets. PINNs approximate unknown field variables using feedforward neural networks to minimize the loss functionals constructed from governing PDEs, initial and boundary conditions at randomly selected collocation points across the domain. PINNs have been successfully applied to solve problems in various domains, including solid mechanics [17, 18], fluid mechanics [19, 20], heat transfer [21, 22], inverse problems [23, 24] to name a few.
In the context of this study, it is crucial to note that, at present, the thermal effects were specifically excluded, directing the attention solely to a 1D coupled hydro-poromechanical analysis. However, it is important to highlight that the framework can be easily extended to model multi-physics processes by incorporating thermal behavior, thus enabling a comprehensive solution for a completely coupled thermo-hydro-poromechanical analysis. Although PINNs were explored as forward solvers to model poroelasticity [25, 26], those were limited to homogeneous materials with only a loose coupling between the governing equations describing solid deformation and fluid flow. The objectives of this study were, therefore, two-fold. Firstly, a modified PINNs architecture capable of modeling poroelastic response for multilayered media characterized by material property discontinuities at interfaces was proposed. Secondly, through several numerical examples in 1D, it was demonstrated that composite neural network architectures are advantageous compared to single neural networks for coupled-field problems. The practical applicability of this framework extends to serving as a numerical solver for interface problems, proving particularly useful in modeling heterogeneous subsurface and geothermal systems with numerous material interfaces due to the presence of geological stratification and numerous faults.
@@
The rest of the paper is organized as follows: Section 2 outlines the governing differential equations for poroelasticity in a heterogeneous domain. Section 3 provides an overview of PINNs, followed by a detailed description of the proposed PINNs architecture tailored for handling poroelasticity in heterogeneous materials. Here, the architecture of the proposed PINNs framework for heterogeneous poroelasticity is described, along with its implementation, and guidelines for improving accuracy and convergence are provided. In Section 4, the numerical results of the proposed PINNs architecture are compared with those obtained using the conventional PINNs approach. The evaluation encompasses numerical accuracy and computational costs on two specific 1D heterogeneous poroelastic problems. Following this analysis, a sensitivity assessment is conducted, wherein the performance of the proposed framework for each applied modification is evaluated. Finally, Section 5 presents concluding remarks and provides an outlook for the work.
@@ 2 Governing Equations
The governing equations have been formulated under the assumptions of quasi-static, linearized poroelasticity [27] such that in a one-dimensional bounded domain $\Omega=[0,L]$, partitioned into two non-overlapping sub-domains $\Omega_{1}=[0,\zeta]$ and $\Omega_{2}=[\zeta,L]$ by a perfectly bonded material interface at $x=\zeta\in(0,L)$ such that $\Omega=\Omega_{1}\cup\Omega_{2}$. The displacements $u_{m}(x,t)$ and pressures $p_{m}(x,t)$ that satisfy the following equations simultaneously in each subdomain $\Omega_{m}$ (for $m=1,2$) are sought:
@@ 3.1 Physics-informed Neural Networks
As described before, the approximation $\tilde{u}(x,t)$ is obtained using a DNN, which is essentially a feed-forward network consisting of a collection of neurons arranged in consecutive layers. Neurons in adjacent layers of the fully connected network are linked, whereas neurons within one single layer are not connected. For a deep neural network with $K$ layers consisting of an input layer, an output layer, and $K-2$ hidden layers, the network approximation can be represented as a combination of $K$ mathematical functions $\mathcal{A}_{i}(\mathbf{x}_{i},\theta_{i})$ as follows:
$\tilde{u}=\mathcal{N}(x,t,\theta)=\mathcal{A}_{K}\circ\mathcal{A}_{K-1}\circ\mathcal{A}_{K-2}\cdot\cdot\cdot\circ\mathcal{A}_{1}(x,t)$ (10)
The layer composition $\circ$ is to be read as $\mathcal{A}_{2}(\mathcal{A}_{1}(\mathbf{x},\theta))$. The function $\mathcal{A}$ can be a linear or non-linear function such that
@@ 3.3 I-PINNs for Coupled PDEs
Regardless of the type of architecture used (SNN or CoNN), the total loss functional aggregates mean square errors (MSEs) from various terms, including governing PDEs (linear momentum and mass balance equations summed together as $\text{MSE}_{\Omega}$), Dirichlet and Neumann boundary conditions for linear momentum balance ($\text{MSE}_{\Gamma_{u}}$ and $\text{MSE}_{\Gamma_{t}}$), and mass balance ($\text{MSE}_{\Gamma_{p}}$, and $\text{MSE}_{\Gamma_{f}}$), the initial conditions for displacements ($\text{MSE}_{IC,u}$), and pressures ($\text{MSE}_{IC,p}$), and the interface compatibility conditions for displacements, tractions, pressures, and fluxes ($\text{MSE}_{\Gamma_{int,u}}$, $\text{MSE}_{\Gamma_{int,\sigma}}$, $\text{MSE}_{\Gamma_{int,p}}$, and $\text{MSE}_{\Gamma_{int,q}}$ respectively). This is represented as:
@@
$\begin{gathered}\text{MSE}_{\Omega}=\frac{1}{N_{\Omega\times T}}\left(\begin{aligned} &\sum_{i=1}^{N_{\Omega\times T}}\left[-\frac{\partial}{\partial x}\left(\nu\frac{\partial\tilde{u}}{\partial x}\right)+\frac{\partial\tilde{p}}{\partial x}\right]\\ &+\sum_{i=1}^{N_{\Omega\times T}}\left[\frac{\partial}{\partial t}\left(a\tilde{p}+\frac{\partial\tilde{u}}{\partial x}\right)-\frac{\partial}{\partial x}\left(\kappa\frac{\partial\tilde{p}}{\partial x}\right)-f_{m}(x,t)\right]\end{aligned}\right)\\[-10.00002pt] \end{gathered}$ (13a)
$\begin{gathered}\text{MSE}_{\Gamma_{u}}=\frac{1}{N_{\Gamma_{u}}}\left(\sum_{i=1}^{N_{\Gamma_{u}}}{\left[\tilde{u}-\overset{\circ}{u_{D}}\right]}\right)\\[-10.00002pt] \end{gathered}$ (13b)
$\begin{gathered}\text{MSE}_{\Gamma_{t}}=\frac{1}{N_{\Gamma_{t}}}\left(\sum_{i=1}^{N_{\Gamma_{t}}}{\left[\nu\frac{\partial\tilde{u}}{\partial x}-\overset{\circ}{u_{N}}\right]}\right)\\[-10.00002pt] \end{gathered}$ (13c)
$\begin{gathered}\text{MSE}_{\Gamma_{p}}=\frac{1}{N_{\Gamma_{p}}}\left(\sum_{i=1}^{N_{\Gamma_{p}}}{\left[\tilde{p}-\overset{\circ}{p_{D}}\right]}\right)\\[-10.00002pt] \end{gathered}$ (13d)
$\begin{gathered}\text{MSE}_{\Gamma_{f}}=\frac{1}{N_{\Gamma_{f}}}\left(\sum_{i=1}^{N_{\Gamma_{f}}}{\left[\kappa\frac{\partial\tilde{p}}{\partial x}-\overset{\circ}{p_{N}}\right]}\right)\\[-10.00002pt] \end{gathered}$ (13e)
$\begin{gathered}\text{MSE}_{\Gamma_{IC}}=\frac{1}{N_{\Gamma_{IC}}}\left(\sum_{i=1}^{N_{\Gamma_{IC}}}{\left[\tilde{u}-\overset{\circ}{u_{IC}}\right]}\right)\\[-10.00002pt] \end{gathered}$ (13f)
$\begin{gathered}\text{MSE}_{\Gamma_{IC}}=\frac{1}{N_{\Gamma_{IC}}}\left(\sum_{i=1}^{N_{\Gamma_{IC}}}{\left[\tilde{p}-\overset{\circ}{p_{IC}}\right]}\right)\\[-10.00002pt] \end{gathered}$ (13g)
$\begin{gathered}\text{MSE}_{\Gamma_{int}}=\frac{1}{N_{\Gamma_{int}}}\left(\sum_{i=1}^{N_{\Gamma_{int}}}{\llbracket\tilde{u}\rrbracket}\right)\\[-10.00002pt] \end{gathered}$ (13h)
$\begin{gathered}\text{MSE}_{\Gamma_{int}}=\frac{1}{N_{\Gamma_{int}}}\left(\sum_{i=1}^{N_{\Gamma_{int}}}{\Biggl{\llbracket}\nu\frac{\partial\tilde{u}}{\partial x}\Biggr{\rrbracket}}\right)\\[-10.00002pt] \end{gathered}$ (13i)
$\begin{gathered}\text{MSE}_{\Gamma_{int}}=\frac{1}{N_{\Gamma_{int}}}\left(\sum_{i=1}^{N_{\Gamma_{int}}}{\llbracket\tilde{p}\rrbracket}\right)\\[-10.00002pt] \end{gathered}$ (13j)

### v3

@@ Abstract
This study presents a novel physics-informed neural network (PINN) framework for modeling poroelasticity in heterogeneous media with material interfaces. The approach introduces a composite neural network (CoNN) where separate neural networks predict displacement and pressure variables for each material. While sharing identical activation functions, these networks are independently trained for all other parameters. To address challenges posed by heterogeneous material interfaces, the CoNN is integrated with the Interface-PINNs or I-PINNs framework (Sarma et al. 2024; https://dx.doi.org/10.1016/j.cma.2024.117135), allowing different activation functions across material interfaces. This ensures accurate approximation of discontinuous solution fields and gradients. Performance and accuracy of this combined architecture were evaluated against the conventional PINNs approach, a single neural network (SNN) architecture, and the eXtended PINNs (XPINNs) framework through two one-dimensional benchmark examples with discontinuous material properties. The results show that the proposed CoNN with I-PINNs architecture achieves an RMSE that is two orders of magnitude better than the conventional PINNs approach and is at least 40 times faster than the SNN framework. Compared to XPINNs, the proposed method achieves an RMSE at least one order of magnitude better and is 40% faster.
@@ 1 Introduction
Coupled-field problems are ubiquitous in the field of subsurface mechanics. Various geothermal and subsurface energy applications are governed by the coupled phenomenon of fluid flow dynamics and solid deformation within poroelastic media. This encompasses activities like hydrocarbon retrieval from oil and gas reservoirs [1], geological storage of carbon dioxide [2], and harnessing energy from subsurface thermal systems [3]. In reservoir engineering [4, 5], these interactions are key in determining the state of subsurface fluids and gases, the performance of production wells, and the deformation of the porous subsurface media. Furthermore, the inherent heterogeneity characterizing these subsurface systems adds a layer of complexity. Heterogeneities, in terms of material interfaces, may stem from the natural stratification of the earth’s crust, the presence of naturally occurring faults, or induced variations due to processes such as hydraulic fracturing [6]. The presence of these interfaces introduces weak or strong discontinuities in these coupled output fields, and their gradients. Therefore, a thorough investigation and comprehension of subsurface heterogeneity in terms of material interfaces and the intricate hydro-poromechanical processes emerge as imperative aspects of studying subsurface applications.
@@
Computational modeling is often necessary to supplement these theoretical and experimental investigations to gain fundamental insight into the coupled physical processes of flow and deformation. These computational models can be broadly categorized as either physics-based or data-driven. Physics-based models involve the solution of coupled partial differential equations (PDEs). These coupled PDEs are solved numerically, often employing methods like the finite element method [7, 8], finite volume method [9], or finite difference method [10], which provide approximate solutions to the PDEs on an underlying grid. However, generating grids that conform to subsurface features (such as material interfaces, faults, and fractures) is a time-consuming pre-processing step. For instance, in several reservoir engineering applications, the cost of grid generation is comparable to (or exceeds) the cost of analysis [11]. This makes computational modeling intractable for realistic field-scale applications. By contrast, data-driven models employ machine learning (ML), particularly artificial neural networks (ANNs), to predict the coupled behavior. Unlike physics-based modeling techniques, these approaches do not rely on an underlying grid for approximation. Instead, they utilize large training datasets and ANNs to predict the physical system’s behavior. However, the precision of these models is highly dependent on the quality and quantity of the training data. This training dataset is often generated through physics-based simulators or experimental methods. Unfortunately, the availability of such training data in subsurface systems is often limited, which restricts the applicability of purely data-driven modeling approaches in poroelasticity. Moreover, data-driven models often perform poorly during extrapolation, i.e. outside the range of training data set.
@@
Physics-informed neural networks (PINNs), a recently developed hybrid approach [12], combines aspects of both physics-based and data-driven modeling. On the one hand, PINNs are not grid-based, thereby eliminating the need for conforming meshes. On the other hand, PINNs leverage fundamental physical laws, including governing PDEs, initial and boundary conditions, and data, to construct a loss functional and thereby do not require extensive training datasets. PINNs approximate unknown field variables using feedforward neural networks to minimize the loss functionals constructed from governing PDEs, initial and boundary conditions at randomly selected collocation points across the domain. PINNs have been successfully applied to solve problems in various domains, including solid mechanics [13, 14], fluid mechanics [15, 16], heat transfer [17, 18], inverse problems [19, 20] to name a few.
Despite these advances, the application of PINNs to poroelasticity has been limited. Notably, Haghighat et al. [21] introduced a sequential training approach based on stress-split algorithms from poromechanics, solving three benchmark homogeneous problems: Mandel’s consolidation problem, Barry–Mercer’s injection-production problem, and a two-phase drainage problem. However, their framework produced poor approximations as the coupling between solid deformation and fluid flow increased. Bekele [22] proposed a PINNs framework to solve Barry and Mercer’s source problem with time-dependent fluid injection/extraction in an idealized poroelastic medium, also in a homogeneous domain. In a follow-up study, Bekele [23] further applied curriculum training to solve a poroelastic problem under idealized conditions. Millevoi et al. [24] addressed the 1D Terzaghi consolidation problem and the 2D Mandel problem, incorporating relatively strong coupling between poroelastic equations, but their approach relied heavily on substantial data extracted from sensors within the domain to support the training process. While these studies provide valuable insights into applying PINNs for poroelasticity, they remain focused on homogeneous systems, often dealing with weakly coupled equations or requiring additional training data to complement the physics-based learning. This highlights a significant research gap in applying PINNs to heterogeneous poroelastic media, especially in scenarios involving strong coupling and material discontinuities.
This paper seeks to bridge this gap by developing an efficient PINNs framework for modeling poroelasticity in heterogeneous media, effectively handling jumps and discontinuities across material interfaces, and managing strong equation coupling based purely on the underlying physics—without the need for additional training data. The specific objectives of this study are threefold:
  • 1.
    To develop a novel framework of PINNs, called composite neural network (CoNN) to address coupled-field problems in poroelasticity with strong fluid-to-solid and solid-to-fluid coupling.
  • 2.
    To integrate the CoNN with the Interface-PINNs (I-PINNs) framework developed by Sarma et al. [25], enabling the model to handle interfacial jumps in field variables, thereby making it suitable for multilayered media with material property discontinuities at interfaces.
  • 3.
    To benchmark the framework, by comparing it against established domain-decomposition-based frameworks available in literature like eXtended PINNs (XPINNs) framework [26], as well as a single neural network architecture.
@@
The practical applications of this framework are significant, particularly as a numerical solver for interface problems. It holds promise for modeling heterogeneous subsurface and geothermal systems characterized by complex geological stratification and numerous material interfaces. The rest of the paper is organized as follows: Section 2 outlines the theory and governing differential equations for poroelasticity in a heterogeneous domain. Section 3 provides an overview of PINNs, followed by a detailed description of the proposed PINNs architecture tailored for handling poroelasticity in heterogeneous materials. Here, the architecture of the proposed PINNs framework for heterogeneous poroelasticity is described, along with its implementation, and guidelines for improving accuracy and convergence are provided. In Section 4, the numerical results of the proposed PINNs architecture are compared with those obtained using the conventional PINNs approach. The evaluation encompasses numerical accuracy and computational costs on two specific 1D heterogeneous poroelastic problems. Following this analysis, a sensitivity assessment is conducted, wherein the performance of the proposed framework for each applied modification is evaluated. Section 5 presents the comparison of the proposed architecture with XPINNs for the set of problems outlined in Section 4. Section 6 presents a discussion of the current study, its limitations, potential for improvements and provides an outlook for the work. Finally, Section 7 presents concluding remarks of this study.
@@ 2 Governing Equations
The governing equations have been formulated under the assumptions of quasi-static, linearized poroelasticity [27] such that in a one-dimensional bounded domain $\Omega=[0,L]$, partitioned into two non-overlapping sub-domains $\Omega_{1}=[0,\zeta]$ and $\Omega_{2}=[\zeta,L]$ by a perfectly bonded material interface at $x=\zeta\in(0,L)$ such that $\Omega=\Omega_{1}\cup\Omega_{2}$. The displacements $u_{m}(x,t)$ and pressures $p_{m}(x,t)$ that satisfy the following equations simultaneously in each subdomain $\Omega_{m}$ (for $m=1,2,...,M$ where $M$ represents the total number of sub-domains) are sought:
@@ 3.1 Physics-informed Neural Networks
As described before, the approximation $\tilde{u}(x,t)$ is obtained using a DNN, which is essentially a feed-forward network consisting of a collection of neurons arranged in consecutive layers. Moreover, the spatiotemporal inputs $(x,t)$ can be collectively wrapped into a single input vector $\mathbf{x}$ such that $\mathbf{x}=[x,t]$. Neurons in adjacent layers of the fully connected network are linked, whereas neurons within one single layer are not connected. For a deep neural network with $K$ layers consisting of an input layer, an output layer, and $K-2$ hidden layers, the network approximation can be represented as a combination of $K$ mathematical functions $\mathcal{A}_{i}(\mathbf{x}_{i},\theta_{i})$ as follows:
$\tilde{u}=\mathcal{N}(\mathbf{x},\theta)=\mathcal{A}_{K}\circ\mathcal{A}_{K-1}\circ\mathcal{A}_{K-2}\cdot\cdot\cdot\circ\mathcal{A}_{1}(\mathbf{x},\theta)$ (10)
The layer composition $\circ$ is to be read as $\mathcal{A}_{i+1}(\mathcal{A}_{i}(\mathbf{x},\theta))$. The function $\mathcal{A}$ can be a linear or non-linear function such that
@@ 3.3 I-PINNs for Coupled PDEs
Regardless of the type of architecture used (SNN or CoNN), the total loss functional $\mathcal{L}(\theta)$ aggregates mean square errors (MSEs) from various terms, given as:
@@
The various loss terms in the above functional, include governing PDEs (linear momentum and mass balance equations summed together as $\text{MSE}_{\Omega}$), Dirichlet and Neumann boundary conditions for linear momentum balance ($\text{MSE}_{\Gamma_{u}}$ and $\text{MSE}_{\Gamma_{t}}$), and mass balance ($\text{MSE}_{\Gamma_{p}}$, and $\text{MSE}_{\Gamma_{f}}$), the initial conditions for displacements ($\text{MSE}_{IC,u}$), and pressures ($\text{MSE}_{IC,p}$), and the interface compatibility conditions for displacements, tractions, pressures, and fluxes ($\text{MSE}_{\Gamma_{int,u}}$, $\text{MSE}_{\Gamma_{int,\sigma}}$, $\text{MSE}_{\Gamma_{int,p}}$, and $\text{MSE}_{\Gamma_{int,q}}$ respectively). These are further represented as:
$\begin{gathered}\text{MSE}_{\Omega}=\frac{1}{N_{\Omega\times T}}\left(\begin{aligned} &\sum_{i=1}^{N_{\Omega\times T}}\left[-\frac{\partial}{\partial x}\left(\nu\frac{\partial\tilde{u}}{\partial x}\right)+\frac{\partial\tilde{p}}{\partial x}\right]\\ &+\sum_{i=1}^{N_{\Omega\times T}}\left[\frac{\partial}{\partial t}\left(a\tilde{p}+\frac{\partial\tilde{u}}{\partial x}\right)-\frac{\partial}{\partial x}\left(\kappa\frac{\partial\tilde{p}}{\partial x}\right)-f_{m}(x,t)\right]\end{aligned}\right)\\ \end{gathered}$ (13a)
$\begin{gathered}\text{MSE}_{\Gamma_{u}}=\frac{1}{N_{\Gamma_{u}}}\left(\sum_{i=1}^{N_{\Gamma_{u}}}{\left[\tilde{u}-\overset{\circ}{u_{D}}\right]}\right)\\ \end{gathered}$ (13b)
$\begin{gathered}\text{MSE}_{\Gamma_{t}}=\frac{1}{N_{\Gamma_{t}}}\left(\sum_{i=1}^{N_{\Gamma_{t}}}{\left[\nu\frac{\partial\tilde{u}}{\partial x}-\overset{\circ}{u_{N}}\right]}\right)\\ \end{gathered}$ (13c)
$\begin{gathered}\text{MSE}_{\Gamma_{p}}=\frac{1}{N_{\Gamma_{p}}}\left(\sum_{i=1}^{N_{\Gamma_{p}}}{\left[\tilde{p}-\overset{\circ}{p_{D}}\right]}\right)\\ \end{gathered}$ (13d)
$\begin{gathered}\text{MSE}_{\Gamma_{f}}=\frac{1}{N_{\Gamma_{f}}}\left(\sum_{i=1}^{N_{\Gamma_{f}}}{\left[\kappa\frac{\partial\tilde{p}}{\partial x}-\overset{\circ}{p_{N}}\right]}\right)\\ \end{gathered}$ (13e)
$\begin{gathered}\text{MSE}_{\Gamma_{IC}}=\frac{1}{N_{\Gamma_{IC}}}\left(\sum_{i=1}^{N_{\Gamma_{IC}}}{\left[\tilde{u}-\overset{\circ}{u_{IC}}\right]}\right)\\ \end{gathered}$ (13f)
$\begin{gathered}\text{MSE}_{\Gamma_{IC}}=\frac{1}{N_{\Gamma_{IC}}}\left(\sum_{i=1}^{N_{\Gamma_{IC}}}{\left[\tilde{p}-\overset{\circ}{p_{IC}}\right]}\right)\\ \end{gathered}$ (13g)
$\begin{gathered}\text{MSE}_{\Gamma_{int}}=\frac{1}{N_{\Gamma_{int}}}\left(\sum_{i=1}^{N_{\Gamma_{int}}}{\llbracket\tilde{u}\rrbracket}\right)\\ \end{gathered}$ (13h)
$\begin{gathered}\text{MSE}_{\Gamma_{int}}=\frac{1}{N_{\Gamma_{int}}}\left(\sum_{i=1}^{N_{\Gamma_{int}}}{\Biggl{\llbracket}\nu\frac{\partial\tilde{u}}{\partial x}\Biggr{\rrbracket}}\right)\\ \end{gathered}$ (13i)
$\begin{gathered}\text{MSE}_{\Gamma_{int}}=\frac{1}{N_{\Gamma_{int}}}\left(\sum_{i=1}^{N_{\Gamma_{int}}}{\llbracket\tilde{p}\rrbracket}\right)\\[-5.0pt] \end{gathered}$ (13j)